OpenStudy (anonymous):

find the integral of 1/sqrt(x^2-9)

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

when you wrote 3cos^2u in this step: "3 sqrt ( 1 - sin^2u) = 3 cos^2 u"... cos is not supposed to be squared right?

OpenStudy (anonymous):

That's why I'm double checking it now! I'm supposedly work out it on paper first.

OpenStudy (anonymous):

= - 9 Int (cosudu/ cosu) = -9 Int ( du/u) = -9 lnu = -9 ln [ arcsin ( x/3) ] + C

OpenStudy (anonymous):

Thanks you for detect the typo --> huge mistake :")

OpenStudy (dominusscholae):

Umm.....sqrt(x^2-9) does not equal -sqrt(9-x^2).....-1 can't be factored out of a square root. Otherwise you would get an imaginary number. Either way, right approach :)

OpenStudy (anonymous):

thank you! could you please explain your steps? i'm a little confused when the u sub is made. also they asked me to integrate it from 3sqrt2 to 6...when the substitutions are made does this change the bounds?

OpenStudy (dominusscholae):

yes. Although I personally do the indefinite integral first and then resbustitute x back in so as to not get confused.

OpenStudy (anonymous):

okay! for my answer i got int (cosu)/(cosu) du ...i do not think this is the correct answer though. after i substitute cos^2(x) in the denom, that is the answer that i got. the coefficients that i took out of the integral was 3/3

OpenStudy (anonymous):

Sorry, my PC and connection in an out so much! Any way, the key is if the Integral is 1/sqrt(9 - x^2), then let x = asinu is correct!

OpenStudy (anonymous):

However in this case, the integrad switch the sign: 1/sqrt(x^2-9)

OpenStudy (anonymous):

That's why I take out negative sign at first to switch it back!

OpenStudy (anonymous):

But after I check the internet, they let x = sec u, then everything just like I work out, expcept with secu.

OpenStudy (anonymous):

Are you following?

OpenStudy (anonymous):

yes, thank you. because i double checked, and i ended up getting du/sqrt(-cosu) which is not possible. i will try it with secu instead

OpenStudy (anonymous):

I'll update with you!

OpenStudy (anonymous):

question: would it be easier to integrate it by parts instead of trig sub?

OpenStudy (anonymous):

No, that's the best method !

OpenStudy (anonymous):

I'll just redo it!

OpenStudy (anonymous):

Let x = 3secu --> dx = 3secu tanu du, x^2 = 9 sec^2u -> sqrt ( x^2 - 9) = sqrt ( 9 sec^2u - 9) = 3 sqrt ( sec^2u -1) = 3 tanu

OpenStudy (anonymous):

Do you understand up to this point?

OpenStudy (anonymous):

yes!

OpenStudy (anonymous):

integral of dx/sqrt(x^2-9) = Int ( 3 sec u tan u du/ 3 tanu ) = Int ( sec u ) du

OpenStudy (anonymous):

ln [ tan u + secu ] + c = ln [ tan ( sec^1 (x/3) + x/3) ] + C

OpenStudy (anonymous):

Do you agree with this result?

OpenStudy (anonymous):

yes, however i don't understand how to find the last step. why is int 1/(secu) = ln [tanu+secu]

OpenStudy (dominusscholae):

sec u =sec u(tan u+ sec u)/(tan u +sec u) = (sec u tan u +sec^2 u)/(tan u +sec u). The numerator is the derivative of the denominator, so using u substitution gets the desured integral.

OpenStudy (anonymous):

oh! okay thank you!